Consider the following investment offers regarding a product you have recently developed. A 10% interest rate should be used throughout this analysis unless otherwise specified:

Offer (I) – Receive $0.55m now and $194k from year 6 through 15. Also, if your product achieved over $100 million in cumulative sales by the end of year 15, you would receive an additional $3m. Assume that there is a 70% probability this would happen.

Offer (II) – Receive 30% of the buyer’s gross profit on the product for the next 4 years. Assume that the buyer’s gross profit margin is 60%. Sales in year 1 are projected to be $2.1m and then expected to grow by 40% per year.

Offer (III) – A trust fund would be set up, calling for semiannual payments of $204k for 8 years. On the 17th period, you would receive the compounded proceeds, which would then be discounted over the 8-year period back to the present at the specified annual rate.

Note: The term “k” is used to represent thousands (× $1,000).

Required: Determine the percentage difference between your most and least profitable alternatives, with the least profitable option as the basis for your calculation.
Answer #1

Offer I:

Present value of $ 194000 receivable from year 6 to 15

Year Amount Disc @ 10% Discounting factor Discounted Cash flows
6 $194,000 ( 1/1.10)^6 0.5645 $109,507.94
7 $194,000 ( 1/1.10)^7 0.5132 $99,552.67
8 $194,000 ( 1/1.10)^8 0.4665 $90,502.43
9 $194,000 ( 1/1.10)^9 0.4241 $82,274.94
10 $194,000 ( 1/1.10)^10 0.3855 $74,795.40
11 $194,000 ( 1/1.10)^11 0.3505 $67,995.82
12 $194,000 ( 1/1.10)^12 0.3186 $61,814.38
13 $194,000 ( 1/1.10)^13 0.2897 $56,194.89
14 $194,000 ( 1/1.10)^14 0.2633 $51,086.26
15 $194,000 ( 1/1.10)^15 0.2394 $46,442.06
Total $740,166.79

Present value of receiving Additional $ 30,00000 if we achieve the Targeted sales = ($ 3 M * 0.70 + $ 0*0.30)/( 1.10)^15

= $ 2100000/( 1.10)^15

= $ 2100000/4.177248

= $ 502723.32

Note: If we will not achieve the Targeted sales we get nothing..

Calculating the present value of amount to be received.

Particulars Amount
Receive $ 0.57 M now $570,000
PV of $ 194000 from year 6 to 15 $740,166.79
PV of $ 30,00000 $502,723.32
Total amount $1,812,890

Hence the Present value of the Future cash inflows is $ 18,12890 under optionI.

Option II:

Year Sales Gross profit ( sales * 60%) Amount ( 30% GP)
1 $2,100,000 $1,260,000.00 $ 1260000*0.30= $ 378000
2 $ 2100000*1.40= $ 2940000 $1,764,000.00 $ 1764000*0.30=$ 529200
3 $ 2940000*1.40=$ 4116000 $2,469,600.00 $ 2469600*0.30=$ 740880
4 $ 4116000*1.40=$ 57,62400 $3,457,440.00 $ 3457440*0.30=$ 10,37232
Total $2,685,312

Computation of the Present value of the Future cash inflows

Year Cash inflow Disc @ 10% Discounted Cash flows
1 $378,000.00 0.9091 $343,636.36
2 $529,200.00 0.8264 $437,355.37
3 $740,880.00 0.7513 $556,634.11
4 $1,037,232.00 0.6830 $708,443.41
Total $2,046,069.26

Hence the Present value of the Future amount is $ 20,46069.26 under option II.

Option III:

Annual Interest rate = 10%

Interest rate for 6 months = 10% /2 = 5%

S.No Amount Future value factor @ 5% Future Value factor Future Cash flows
1

$204000

(1.05)^15 2.0789 $424101.35
2 $204000 ( 1.05)^14 1.9799 $403906.05
3 $204000 ( 1.05)^13 1.8856 $384672.43
4

$204000

( 1.05)^12 1.7959 $366354.69
5 $204000 ( 1.05)^11 1.7103 $348909.23
6 $204000 ( 1.05)^10 1.6289 $332294.5
7 $204000 ( 1.05)^9 1.5513 $316470.96
8 $204000 ( 1.05)^8 1.4775 $301400.91
9 $204000 ( 1.05)^7 1.4071 $287048.49
10 $204000 ( 1.05)^6 1.3401 $ 273379.51
11 $204000 ( 1.05)^5 1.2763 $ 260361.44
12 $204000 ( 1.05)^4 1.2155 $247963.28
13 $204000 ( 1.05)^3 1.1576 $236155.50
14 $204000 ( 1.05)^2 1.1025 $ 224910
15 $204000 ( 1.05)^1 1.0500 $214200
16 $204000 ( 1.05)^0 1.0000 $204000
Total $4826128.32

It is assumed that funds are reinvested after every cash flow

It is assumed that semi Annual payments occur at the end of each year.

Present value of $ 4826128.32is $ 4826128.32/( 1.10)^8

=$ 4826128.32/( 1.10)^8

= $ 2251424.479

Present value under option III is $ 2251424.79

Particulars Amount Status
Option 1 Present value $1,812,890 Least Profitable Alternative
Option II Present value $2,046,069.26
Option III Present value $2251424.79 Most profitable Alternative

% Difference between Most profitable Alternative and least profitable alternative is ( $ 2251424.79-$ 1812890)/$ 1812890

= $ 438534.79/$ 1812890*100

= 24.1898%

Hence there is 24.1898% diffeerence between most Profitable and least profitable alternative.